Show that the function f: R^2→R^2 given by

f(x,y) = (xy^3+1, x^2+y^2) is not invertible. Futher

check whether it is locally invertible at the point (2,1)



1
Expert's answer
2021-05-25T15:27:20-0400

According to the Inverse function theorem, a function is invertible in a neighborhood of a point in its domain: namely, that its Jacobian is continuous and non-zero at the point. 

Let us build the Jacobian:

"J=
\begin{pmatrix}
 f_{1x} & f_{1y} \\
 f_{2x} & f_{2y}
\end{pmatrix} =
\begin{pmatrix}
 y^3 & 3xy^2 \\
 2x & 2y
\end{pmatrix}"

and the Jacobian determinant is


"\det J=2y^4-6x^2y^2=2y^2(y^2-3x^2)"

"\det J=0 \iff y=0;\quad or \quad y=\pm\sqrt{3}x"

Therefore this function is not invertible.

At the point (2,1)

"\det J(2,1)=2*1^2(1^2-3*2^2)=
2∗1 
2
 (1 
2
 −3∗2 
2
 )=-22\ne 0"

Therefore this function is locally invertible at (2,1)